OpenStudy (anonymous):

36 to the power of what = 1000? And how do I work that out? Will be happy to pay 5 bucks on paypal if you can help me. Thanks

OpenStudy (anonymous):

36^x = 1000 take log of both sides.

OpenStudy (anonymous):

How to do that? My math is not very good

OpenStudy (anonymous):

Have you studied logarithms?

OpenStudy (anonymous):

Never, I am writing a computer program and need to use this to select numbers.

OpenStudy (anonymous):

well........if this is not for a math class.......just for your program...I can tell you that the answer is a decimal...approximately 1.927645813

OpenStudy (anonymous):

Yes it is for my program but I need to do this for many numbers. So that is why I need to work out how. Is it complex?

OpenStudy (anonymous):

This decimal is correct to the nearest ten- millionth.

OpenStudy (anonymous):

You would need to take the log of both sides of the equation, you get x(log 36) = log (1000) Divide both sides by log 36, you get x = log(1000)/log(36)= 1.927645813 We are using the Law of Logs that states that log(a^r) = r(log a) So, log (36^x) = x(log 36)

OpenStudy (anonymous):

Ok thank you can you help me with these problems I am happy to pay. Do you have an email?

OpenStudy (anonymous):

No pay. Please. We are here to help. If you were a student, I would expect you to work along with me using your knowledge of logs. But you are using this for a program, mathematically related, so we helped you out. Take care.

OpenStudy (anonymous):

Thanks you are very nice, have a good day to you too.

OpenStudy (anonymous):

Welcome.